Gail Dribble is analyzing the shares of Petscan Radiology. Petscan’s shares pay a dividend once each year, and it just distributed this year’s $0.85 dividend. The market price of each of the shares is $12.14. Gail estimates that Petscan will increase its dividend by 7% per year forever. After contemplating the risk of Petscan shares, Gail is willing to hold the shares only if they provide an annual expected return of at least 13%. Should she buy Petscan shares or not?
Answer #1

let us estimate the fair price of the share using dividend growth model:

P = d(1+g) / (k-g)

here,

d = $0.85

g=7%=>0.07

k = 13%

=>0.13

price = $0.85*(1.07) / (0.13 - 0.07)

=>0.9095/0.06

=>$15.16.

since the actual price of $12.14 is less than the fair price of $15.16 Gail driblle can buy the shares.

or

Alternate presentation:

let us know the required return offered by the stock at a price of $12.14

=> price = d(1+g) / (k-g)

here k is to be found out

=>12.14 = $0.85*(1.07) / (k-0.07)

=>k-0.07 = 0.9095 /12.14

=>k-0.07 = 0.0749176

=>k=0.1449 or 14.49%

since the return offered is greater than the expected return of 13%, the share can be bought by Gail Dribble.